OpenStudy (anonymous):

To what number is 29 connected by a whisker in this data set?

OpenStudy (anonymous):

10, 11, 13, 15, 17, 18, 20, 22, 23, 25, 27, 29

OpenStudy (anonymous):

Ok. To calculate the number, we start with 10.

OpenStudy (anonymous):

We take the difference of each pair of numbers.

OpenStudy (anonymous):

11 - 10 = 1 13 - 11 = 2 15 - 13 = 2 ... and so on until 29 - 27 = 2

OpenStudy (anonymous):

The answer is 10 + {the sum of difference of each pair}

OpenStudy (campbell_st):

in the data set you need the 5 number summary... the high is 29 and it is connected by a whisker to the upper quartile or 3rd quartile. the median is 19 so find the middle of the upper half of the data, that will be the upper quartile and the numbe rconnected to 29